# Closed-Form Feedback-Free Learning with Forward Projection

### v1

@@ Results
For backpropagation, $\mathcal{O}(m^{2})$ memory is required for storing layer weights, $\mathcal{O}(m^{2})$ for accumulated gradients, and $\mathcal{O}(m)$ for activations. Here, the computation requirements scale as $\mathcal{O}(N_{e}m^{2})$ MACs each for the forward pass, backward pass, and weight update calculations. Hence, in the typical setting where $m>>m_{L}$, we note that forward projection and backpropagation have similar memory requirements, scaling as $\mathcal{O}(m^{2})$. Notably, compute scales as $\mathcal{O}(m^{3})$ for FP versus $\mathcal{O}(N_{e}m^{2})$ for BP. For example, training a dense hidden layer with $1000$ inputs and $1000$ outputs on the FMNIST dataset ($N=60,000,m_{L}=10$) by forward projection requires a total of $1.2\times 10^{11}$ MAC operations for the forward pass and $1.2\times 10^{11}$ operations for the weight update. Training the same layer via 100 epochs of backpropagation requires $6.0\times 10^{12}$ MAC operations for forward passes and $1.2\times 10^{13}$ operations for weight updates. Forward projection requires only a single training epoch, leading to a proportional speed up in the overall computation time. Note that local supervision and forward-forward algorithms have similar computational and memory requirements as backpropagation, as illustrated in Table 2. We also present the wall-clock time for full training of an MLP with $3\times 1000$ hidden neurons on FMNIST, demonstrating a 500×500$\times$500 × speedup with forward projection in this example.
Compute Memory Time
Method Training epochs Forward pass Weight update Model parameters Weight update Training time (s)
Backprop-agation $N_{e}$ $N_{e}Nm^{2}$ $2N_{e}Nm^{2}$ $m^{2}$ $m^{2}+2m$ $248.8\pm 2.9$
Local Supervision $N_{e}$ $N_{e}N(m^{2}+mm_{L})$ $N_{e}N(m^{2}+mm_{L})$ $m^{2}+mm_{L}$ $m^{2}+2m$ $472.0\pm 5.6$
Forward-Forward $N_{e}$ $2N_{e}Nm^{2}$ $N_{e}Nm^{2}$ $m^{2}$ $m^{2}$ $648.2\pm 3.8$
Forward Proj. (Ours) $1$ $N(2m^{2}+mm_{L})$ $2Nm^{2}+2m^{3}$ $2m^{2}+mm_{L}$ $2m^{2}$ ${0.44\pm 0.1}$
Table 2: Complexities of each training method in an MLP with $m$ inputs, one output and $L$ hidden layers of dimension $m$. $N$: training sample size. $N_{e}$: epochs, all parameters except activations are per layer.
@@ Discussion
We present a novel algorithm, Forward Projection, which enables learning in a single pass over the dataset using random projections and closed-form optimisation. Compared to state-of-the-art local learning methods that require observing post-synaptic neuronal outputs to optimise synaptic weights via error-based gradient descent, Forward Projection operates under a stricter constraint, fitting weights using just the pre-synaptic neuronal activity and labels.
@@
Few-shot learning is ubiquitous in biological systems, which exhibit rapid neuronal adaptation to changing environments [32]. In our experiments, Forward Projection demonstrated clear performance advantages over backpropagation and local learning approaches in few-shot learning tasks, presenting a plausible method for learning new tasks rapidly. In this setting, convolutional features learned by Forward Projection yielded more generalisable models than backpropagation, which overfitted in some experiments and underfitted in others, even underperforming random features in some cases. Forward Projection performed similarly to local supervision and forward-forward training in these tasks. FP also maintained reasonable performance on activation functions which were untrainable by SGD-based methods (see section A.7 in SI).
The closed-form Forward Projection fit is computable in a single pass over the data for each layer, presenting a novel opportunity to expedite training and reduce environmental footprint. The efficiency of FP training is attributable to a substantially different operational sequence to that employed in SGD-based approaches. Firstly, FP training collects a gram matrix of pre-synaptic activity over a single epoch, which is subsequently inverted during a one-step weight matrix computation. Secondly, FP completes fitting for each layer before initialising successive layers. In contrast, SGD-based approaches fit all the layers in the network iteratively, using feedback from neuronal outputs and downstream layers. FP training requires no retrograde communication between neural output activations and pre-synaptic connections, enabling direct training on hardware with unidirectional synaptic and neuronal communication. This differs from current iterative local learning methods, which require backward communication from neuronal outputs to optimise pre-synaptic parameters (see section A.1 in SI). Although FP addresses the feedback-free constraint of biological learning systems, the biological plausibility of one-step learning through matrix inversion is uncertain. Forward Projection training may be applicable as a pre-training step to reduce the number of training epochs required for backpropagation.
In conclusion, FP is a novel approach for neural network optimisation, employing techniques from randomised projective embedding and linear regression to fit weight matrices in one epoch using a single-step solution. Interpretability of hidden neurons in FP-trained models may be employed to improve the explainability of neural network predictions.
@@ Methods
The PTB-XL dataset contains 12-lead electrocardiography (ECG) recordings from 18,889 participants [25, 34]. ECG recordings of ten-second duration and 100Hz sample rate were used in our experiments. The predictive task was to discriminate ECG recordings with normal waveform morphology ($N_{train}=6,451;N_{test}=721$) from those diagnosed as myocardial infarction ($N_{train}=2,707;N_{test}=268$) by a cardiologist. Data instances with uncertain diagnoses were excluded from this analysis. Following recommendations of the data providers who provided predefined participant-disjoint dataset splits, the tenth fold was held out for model testing [25].
@@
The Human Non-TATA Promoters dataset (“Promoters”) was extracted from the GenomicBenchmarks repository [24]. Data was originally published in [23]. 36,131 nucleotide sequences of 251 bases each were analysed. Nucleotide sequences were converted to 4-channel one-hot vectors indicating adenine, cytosine, guanine and thymine. Indeterminate bases were represented with zero vectors. Models were required to classify the promoter functionality of the sequence as “promoter” ($N_{train}=12,355;N_{test}=4,119$) or “non-promoter” ($N_{train}=14,742;N_{test}=4,915$).
@@
The paediatric pneumonia chest X-ray dataset (“CXR”) is a retrospective cohort of patients aged between one and five years, recorded in Guangzhou Women and Children’s Medical Center, Guangzhou, originally published in [27]. Chest X-ray images were recorded as part of routine care during diagnostic workup for suspected lower respiratory tract infection. During data collection, clinicians screened the images for quality and excluded those with severe artefact or corruption. Images were annotated by two expert physicians. Local institutional review board approvals were obtained. Images were loaded in greyscale, rescaled to the $[0,1]$ intensity range and resized to $128\times 128$ pixels by bilinear interpolation. Models were required to classify images “normal” ($N_{train}=1,349;N_{test}=234$), “viral pneumonia” ($N_{train}=2,538;N_{test}=242$), or “bacterial pneumonia” ($N_{train}=1,345;N_{test}=148$).
@@
The optical coherence tomography dataset (“OCT”) is a retrospective cohort of adult patients from five ophthalmology institutions in the USA and China recorded between 2013 and 2017 during routine care, originally published in [27]. Images were initially annotated by local medical students, who had received OCT interpretation training. Subsequent annotation was performed by four ophthalmologists and two independent retinal specialists. Horizontal foveal cut images were available in portable network graphics image format. Images were loaded in greyscale, rescaled to the $[0,1]$ intensity range, and resized to $128\times 128$ pixels by bilinear interpolation. Models were required to classify images as one of “normal” ($N_{train}=2,926;N_{test}=149$) or “choroid neovascularisation” ($N_{train}=791;N_{test}=178$).
@@
The FMNIST dataset was modelled using an MLP with $3\times 1000$ hidden ReLU-activated neurons. Sequential datasets (PTBXL-MI and Promoters) were modelled by a 1d-CNN architecture of four convolutional blocks. Each convolutional block included two convolutional layers with kernel dimension $3$, and strides of 1 and 2 respectively. Convolutional layers in the $l$-th block had $32\times 2^{l-1}$ filters. Convolutional implementation is detailed in Supplementary section A.2. Convolutional outputs were aggregated by global average pooling in the penultimate layer. For gradient-descent-based learning algorithms, batch normalisation layers were included between convolutional blocks. Image datasets (CXR and OCT) were modelled by few-shot learning using an equivalent 2d-CNN architecture with kernel dimension $3\times 3$ and $16\times 2^{l-1}$ filters in the $l$-th convolutional block. The $\operatorname{sign}$ function was employed to generate target activations for forward projection models, such that
@@
Data augmentation was not performed in our experiments. Models were fitted using the PyTorch library. For SGD-based methods, early stopping was performed according to validation loss. Model weight initialisation was random, and optimisation was performed with the Adam optimiser with a learning rate of 0.001, training to minimise validation loss by early stopping with a patience of five epochs. In few-shot experiments, the learning rate was reduced to 0.0001 for SGD-based models and a patience of ten epochs was employed. Models were fitted to minimise categorical or binary cross-entropy as appropriate. Forward-forward and local supervision implementation are detailed in supplementary section A.3.
@@
The authors declare no competing interests.
@@
The PTBXL-MI dataset is available from PhysioNet [34]. The Promoters dataset is available from the GenomicBenchmarks library [24] and the original publication [23]. The CXR and OCT datasets are provided in [27].
@@ A.2 Convolutional Layer Implementation for FP
$\tilde{\textbf{z}}_{l,r,c}=g_{l}(\textbf{a}_{l-1,r,c}\textbf{Q}_{l})+g_{l}(\textbf{yU}_{l}).$ (6)
@@
$\textbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{A}_{l-1,r,c}^{T}\textbf{A}_{l-1,r,c}+\lambda I\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{A}_{l-1,r,c}^{T}\tilde{\textbf{Z}}_{l,r,c}\right).$ (7)
Here, $\textbf{A}_{l-1,r,c}\in\mathbb{R}^{N\times(m_{l-1}k_{1}k_{2})}$ is a matrix representing the pre-synaptic input for the convolutional kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. Likewise, $\tilde{\textbf{Z}}_{l,r,c}\in\mathbb{R}^{N\times m_{l}}$ contains target potentials for the kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. In each task, the FP penalty parameter was fixed at $\lambda=10$ for hidden layers and $\lambda=1$ for output layers.
A.3 Implementation of Local Supervision and Forward Forward

### v2

@@ Results
For backpropagation, $\mathcal{O}(m^{2})$ memory is required for storing layer weights, $\mathcal{O}(m^{2})$ for accumulated gradients, and $\mathcal{O}(m)$ for activations. Here, the computation requirements scale as $\mathcal{O}(N_{e}m^{2})$ MACs each for the forward pass, backward pass, and weight update calculations. Hence, in the typical setting where $m>>m_{L}$, we note that Forward Projection and backpropagation have similar memory requirements, scaling as $\mathcal{O}(m^{2})$. Notably, compute scales as $\mathcal{O}(m^{3})$ for FP versus $\mathcal{O}(N_{e}m^{2})$ for BP. For example, training a dense hidden layer with $1000$ inputs and $1000$ outputs on the FMNIST dataset ($N=60,000,m_{L}=10$) by Forward Projection requires a total of $1.2\times 10^{11}$ MAC operations for the forward pass and $1.2\times 10^{11}$ operations for the weight update. Training the same layer via 100 epochs of backpropagation requires $6.0\times 10^{12}$ MAC operations for forward passes and $1.2\times 10^{13}$ operations for weight updates. Forward projection requires only a single training epoch, leading to a proportional speed up in the overall computation time. Note that Local Supervision and Forward-Forward algorithms have similar computational and memory requirements as backpropagation, as illustrated in Table 2. We also present the wall-clock time for full training of an MLP with $3\times 1000$ hidden neurons on FMNIST, demonstrating a $66\times$ speedup with Forward Projection in this example. Further timing results are provided in Table5 in SI. Computations were run using the Google Colab service with an NVIDIA T4 graphics processing unit.
Compute Memory FMNIST Training
Method Training epochs Forward pass Weight update Model parameters Weight update Time (s) Epochs
Backprop-agation $N_{e}$ $N_{e}Nm^{2}$ $2N_{e}Nm^{2}$ $m^{2}$ $m^{2}+2m$ $22.43\pm 4.8$ $11.6\pm 2.7$
Local Supervision $N_{e}$ $N_{e}N(m^{2}+mm_{L})$ $N_{e}N(m^{2}+mm_{L})$ $m^{2}+mm_{L}$ $m^{2}+2m$ $46.9\pm 12.5$ $10.8\pm 3.56$
Forward-Forward $N_{e}$ $2N_{e}Nm^{2}$ $N_{e}Nm^{2}$ $m^{2}$ $m^{2}$ $172\pm 56$ $40.2\pm 13.8$
Predictive Coding $N_{e}$ $N_{e}Nm^{2}$ $4N_{e}Nm^{2}$ $2m^{2}$ $2m^{2}+2m$ $150\pm 61$ $25.4\pm 10.2$
Difference Target Prop. $N_{e}$ $N_{e}Nm^{2}$ $8N_{e}Nm^{2}$ $2m^{2}$ $2m^{2}+4m$ $72.7\pm 26.4$ $13.6\pm 5.3$
Forward Proj. (Ours) 11 $N(2m^{2}+mm_{L})$ $2Nm^{2}+2m^{3}$ $2m^{2}+mm_{L}$ $2m^{2}$ $0.34\pm 0.1$} 11
Table 2: Complexities of each training method in an MLP with $m$ inputs, one output and $L$ hidden layers of dimension $m$. NN: training sample size. $N_{e}$: epochs, all parameters except activations are per layer.
@@ Discussion
We present the Forward Projection algorithm, which enables learning in a single pass over the dataset using random projections and closed-form optimisation. Compared to state-of-the-art local learning methods that require observing post-synaptic neuronal outputs to optimise synaptic weights via error-based gradient descent, Forward Projection operates under a stricter constraint, fitting weights using just the pre-synaptic neuronal activity and labels.
@@
Few-shot learning is ubiquitous in biological systems, which exhibit rapid neuronal adaptation to changing environments [35]. In our experiments, Forward Projection demonstrated clear performance advantages over backpropagation and local learning approaches in few-shot learning tasks, presenting a plausible method for learning new tasks rapidly. In this setting, convolutional features learned by Forward Projection yielded more generalisable models than backpropagation, which overfitted in some experiments and underfitted in others, even underperforming random features in some cases. Forward Projection performed similarly to Local Supervision and Forward-Forward training in these tasks. Forward Projection also maintained reasonable performance on activation functions which were untrainable by SGD-based methods (see section A.11 in SI).
The closed-form Forward Projection fit is computable in a single pass over the data for each layer, presenting an opportunity to expedite training and reduce environmental footprint. The efficiency of FP training is attributable to a substantially different operational sequence to that employed in SGD-based approaches. Firstly, FP training collects a gram matrix of pre-synaptic activity over a single epoch, which is subsequently inverted during a one-step weight matrix computation. Secondly, FP completes fitting for each layer before initialising successive layers. In contrast, SGD-based approaches fit all the layers in the network iteratively, using feedback from neuronal outputs and downstream layers. FP training requires no retrograde communication between neural output activations and pre-synaptic connections, enabling direct training on hardware with unidirectional synaptic and neuronal communication. This differs from current iterative local learning methods, which require backward communication from neuronal outputs to optimise pre-synaptic parameters (see section A.1 in SI). Although FP addresses the feedback-free constraint of biological learning systems, the biological plausibility of one-step learning through matrix inversion is uncertain. Forward Projection training may be applicable as a pre-training step to reduce the number of training epochs required for backpropagation.
In conclusion, FP is an efficient approach for neural network optimisation, employing techniques from randomised projective embedding and linear regression to fit weight matrices in one epoch using a single-step solution. Interpretability of hidden neurons in FP-trained models may be employed to improve the explainability of neural network predictions.
@@ Methods
The PTB-XL dataset contains 12-lead electrocardiography (ECG) recordings from 18,889 participants [28, 37]. ECG recordings of ten-second duration and 100Hz sample rate were used in our experiments. The predictive task was to discriminate ECG recordings with normal waveform morphology ($N_{\text{train}}=6,451;N_{\text{test}}=721$) from those diagnosed as myocardial infarction ($N_{\text{train}}=2,707;N_{\text{test}}=268$) by a cardiologist. Data instances with uncertain diagnoses were excluded from this analysis. Following recommendations of the data providers who provided predefined participant-disjoint dataset splits, the tenth fold was held out for model testing [28]. The first 15 MI-positive ECGs in the test set were annotated by a medical doctor to identify diagnostically relevant sections for quantitative explainability evaluation.
@@
The Human Non-TATA Promoters dataset (“Promoters”) was extracted from the GenomicBenchmarks repository [27]. Data was originally published in [26]. 36,131 nucleotide sequences of 251 bases each were analysed. Nucleotide sequences were converted to 4-channel one-hot vectors indicating adenine, cytosine, guanine and thymine. Indeterminate bases were represented with zero vectors. Models were required to classify the promoter functionality of the sequence as “promoter” ($N_{\text{train}}=12,355;N_{\text{test}}=4,119$) or “non-promoter” ($N_{\text{train}}=14,742;N_{\text{test}}=4,915$).
@@
The paediatric pneumonia chest X-ray dataset (“CXR”) is a retrospective cohort of patients aged between one and five years, recorded in Guangzhou Women and Children’s Medical Center, Guangzhou, originally published in [30]. Chest X-ray images were recorded as part of routine care during diagnostic workup for suspected lower respiratory tract infection. During data collection, clinicians screened the images for quality and excluded those with severe artefact or corruption. Images were annotated by two expert physicians. Local institutional review board approvals were obtained. Images were loaded in greyscale, rescaled to the $[0,1]$ intensity range and resized to $128\times 128$ pixels by bilinear interpolation. Models were required to classify images “normal” ($N_{\text{train}}=1,349;N_{\text{test}}=234$), “viral pneumonia” ($N_{\text{train}}=2,538;N_{\text{test}}=242$), or “bacterial pneumonia” ($N_{\text{train}}=1,345;N_{\text{test}}=148$).
@@
The optical coherence tomography dataset (“OCT”) is a retrospective cohort of adult patients from five ophthalmology institutions in the USA and China recorded between 2013 and 2017 during routine care, originally published in [30]. Images were initially annotated by local medical students, who had received OCT interpretation training. Subsequent annotation was performed by four ophthalmologists and two independent retinal specialists. Horizontal foveal cut images were available in portable network graphics image format. Images were loaded in greyscale, rescaled to the $[0,1]$ intensity range, and resized to $128\times 128$ pixels by bilinear interpolation. Models were required to classify images as one of “normal” ($N_{\text{train}}=2,926;N_{\text{test}}=149$) or “choroid neovascularisation” ($N_{\text{train}}=791;N_{\text{test}}=178$).
@@
The FMNIST dataset was modelled using an MLP with $3\times 1000$ hidden ReLU-activated neurons. Sequential datasets (PTBXL-MI and Promoters) were modelled by a 1d-CNN architecture of four convolutional blocks. Each convolutional block included two convolutional layers with kernel dimension $3$, and strides of 1 and 2 respectively. Convolutional layers in the $l$-th block had $32\times 2^{l-1}$ filters. Convolutional implementation is detailed in Supplementary section A.2. Convolutional outputs were aggregated by global average pooling in the penultimate layer. For gradient-descent-based learning algorithms, batch normalisation layers were included between convolutional blocks. CIFAR modelling in Table 1 used a vision transformer architecture [38] operating on image patches of dimension $4\times 4$, with a sequential stack of four multi-headed attention layers, each having 8 heads, embedding dimension 64, and MLP dimension 64. Few-shot image modelling tasks used a 2d-CNN architecture with kernel dimension $3\times 3$ and either $16\times 2^{l-1}$ filters in the $l$-th convolutional block (CXR and OCT) or $32\times 2^{l-1}$ filters in the $l$-th convolutional block (CIFAR). The $\operatorname{sign}$ function was employed to generate target activations for Forward Projection models, such that
@@
Data augmentation was not performed in our experiments. Models were fitted using the PyTorch library. For SGD-based methods, early stopping was performed according to validation loss. Model weight initialisation was random, and optimisation was performed with the Adam optimiser with a learning rate of 0.001, training to minimise validation loss by early stopping with a patience of five epochs. In few-shot experiments, the learning rate was reduced to 0.0001 for SGD-based models and a patience of ten epochs was employed. It is acknowledged that direct comparison of gradient descent based training with closed-form solutions is not strictly “like-for-like”. However, all efforts were made to ensure experimental conditions were otherwise equivalent. Models were fitted to minimise categorical or binary cross-entropy as appropriate. Implementations of Forward-Forward, Local Supervision, Difference Target Propagation and Predictive Coding methods are detailed in supplementary section A.4.
@@
Robert O’Shea contributed to the initial draft of the manuscript whilst employed at King’s College London. He commenced employment with Eli Lilly & Company prior to the submission of the revised manuscript. His involvement in this project is entirely independent of his role at Eli Lilly & Company and was not influenced by his employment there.
@@
The PTBXL-MI dataset is available from PhysioNet [37]. The Promoters dataset is available from the GenomicBenchmarks library [27] and the original publication [26]. The CXR and OCT datasets are provided in [30]. Pre-processed datasets and cross-validation folds are available for download at fp_datasets - Mendeley Data [39].
@@ A.2 Convolutional Layer Implementation for FP
$\tilde{\textbf{z}}_{l,r,c}=g_{l}\left(\textbf{a}_{l-1,r,c}\textbf{Q}_{l}\right)+g_{l}\left(\textbf{yU}_{l}\right).$ (6)
@@
$\textbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{A}_{l-1,r,c}^{T}\textbf{A}_{l-1,r,c}+\lambda\textbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\textbf{A}_{l-1,r,c}^{T}\tilde{\textbf{Z}}_{l,r,c}\right).$ (7)
Here, $\textbf{A}_{l-1,r,c}\in\mathbb{R}^{N\times(m_{l-1}k_{1}k_{2})}$ is a matrix representing the pre-synaptic input for the convolutional kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. Likewise, $\tilde{\textbf{Z}}_{l,r,c}\in\mathbb{R}^{N\times m_{l}}$ contains target potentials for the kernel in row $r$ and column $c$ of the $l$-th layer over all $N$ training instances. In each task, the FP penalty parameter was fixed at $\lambda=10$ for hidden layers and $\lambda=1$ for output layers. To prevent numerical overflow due to large layers or training datasets, a scalar reduction factor $\tau\in(0,1]$may be applied to downscale $\textbf{A}^{T}\textbf{A}$, $\lambda$, and $\textbf{A}^{T}\textbf{Z}$ as necessary, such that:
$\textbf{W}_{l}:=\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\tau\textbf{A}_{l-1,r,c}^{T}\textbf{A}_{l-1,r,c}+\tau\lambda\textbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\sum^{C}_{c=1}\tau\textbf{A}_{l-1,r,c}^{T}\tilde{\textbf{Z}}_{l,r,c}\right).$ (8)
A.3 Attention Layer Implementation for FP
We consider the implementation of Forward Projection in a transformer layer over pre-synaptic input $\textbf{a}_{l-1}\in\mathbb{R}^{1\times R\times m_{l-1}}$, with $R$ patches, each having $m_{l-1}$ channels. Each attention head of dimension $d\in\mathbb{N}$ contains three weight matrices for optimisation: $\textbf{W}_{l}^{\operatorname{Query}}$, $\textbf{W}_{l}^{\operatorname{Key}}$ and $\textbf{W}_{l}^{\operatorname{Value}}$, yielding intermediate activations $\textbf{z}_{l}^{\operatorname{Query}}=\textbf{a}_{l-1}\textbf{W}_{l}^{\operatorname{Query}}$, $\textbf{z}_{l}^{\operatorname{Key}}=\textbf{a}_{l-1}\textbf{W}_{l}^{\operatorname{Key}}$, and $\textbf{z}_{l}^{\operatorname{Value}}=\textbf{a}_{l-1}\textbf{W}_{l}^{\operatorname{Value}}$ respectively. The attention scores $\textbf{s}_{l}$ are then calculated as
$\textbf{s}_{l}:={\operatorname{Softmax}\left({\frac{\textbf{z}_{l}^{\operatorname{Query}}\times(\textbf{z}_{l}^{\operatorname{Key}})^{T}}{\sqrt{d}}}\right)\times\textbf{z}_{l}^{\operatorname{Value}}}$ (9)
Forward projection may be applied to generate target values for each of $\textbf{z}_{l}^{\operatorname{Query}},\textbf{z}_{l}^{\operatorname{Key}},\textbf{z}_{l}^{\operatorname{Value}}$, using distinct data projection matrices $\textbf{Q}_{l}^{\operatorname{Query}},\textbf{Q}_{l}^{\operatorname{Key}},\textbf{Q}_{l}^{\operatorname{Value}}$, and a shared label projection matrix $\textbf{U}_{l}$, such that for each patch, query matrix target potentials are given by
$\tilde{\textbf{z}}_{l,r}^{\operatorname{Query}}=g_{l}\left(\textbf{a}_{l-1,r}\textbf{Q}_{l}^{\operatorname{Query}}\right)+g_{l}\left(\textbf{yU}_{l}\right).$ (10)
Thus, the query weight matrix is given by the penalised regression solution. Weights were fitted such that
$\textbf{W}_{l}^{\operatorname{Query}}:=\left(\sum^{R}_{r=1}\textbf{A}_{l-1,r}^{T}\textbf{A}_{l-1,r}+\lambda\textbf{I}\right)^{-1}\left(\sum^{R}_{r=1}\textbf{A}_{l-1,r}^{T}\tilde{\textbf{Z}}_{l,r}^{\operatorname{Query}}\right).$ (11)
$\textbf{W}_{l}^{\operatorname{Key}}$ and $\textbf{W}_{l}^{\operatorname{Value}}$, are fitted in the same manner - using $\textbf{Q}_{l}^{\operatorname{Key}}$ and $\textbf{Q}_{l}^{\operatorname{Value}}$, respectively, in place of $\textbf{Q}_{l}^{\operatorname{Query}}$ in (10).
A.4 Implementation of local learning methods